identify the amount,base,and p value of 25% of 400cars is 100 cars
45 of 50 children is 90%
in a recent survey38% ofUnited States say that gas prices have affected the type 0f vehecle they will buy. In a sample of 500 people who are in the market for a new vehcle,how many would you expect to be influenced by gas prices?
Yellowstone National Park has 3366mi.2 of undevelopedland. If this represents 99% of the totalarea,fine the total area of the park?
As of the2002--2003 football season,Peyton Manning had completed 1357 passes out of 2226 attempts. Find his completion percetage to the nearest tenth of a percent.

Answer by stanbon(75887)

identify the amount,base,and p value of 25% of 400cars is 100 cars
amount: 100
base: 400
p-value: 25 or maybe 25% (depends on what your text says)
-------------------------------
45 of 50 children is 90%
mount: 45
base: 50
p-value: 90 or 90%
-------------------------------
in a recent survey38% ofUnited States say that gas prices have affected the type 0f vehecle they will buy. In a sample of 500 people who are in the market for a new vehcle,how many would you expect to be influenced by gas prices?
---
18% of 500 = 90 people
-------------------
Yellowstone National Park has 3366mi.2 of undevelopedland. If this represents 99% of the total area, ,find the total area of the park?
---
0.99*(total area) = 3366 sq mi
total area = 3366/0.99 = 3400 sq. mi.
--------------------------------
As of the 2002--2003 football season, Peyton Manning had completed 1357 passes out of 2226 attempts. Find his completion percentage to the nearest tenth of a percent.
1357/2226 = 0.60961 = 61%
================================
Cheers,
Stan H.